# TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling

### v1

@@ 2.4 TriangleMix
[…]
Table 2: Comparison with static sparsity methods on Longbench. TriangleMix is our proposed method. StreamingMix is a hybrid baseline that uses dense attention in shallow layers but replaces Triangle attention with a streaming pattern in deeper layers. The number of dense layers matches that of TriangleMix.
From Figure 4, we observe that both the Middle Q-K and Last Q-K sections have low average attention scores, with the Middle Q-K section slightly higher across all layers. However, when examining the gradient values, we find that the derivatives in the Last Q-K section are significantly larger than those in the Middle Q-K section. This indicates that the model output is more sensitive to changes in the attention scores of the Last Q-K section compared to the Middle Q-K section.
Besides, for the last layer, we observe that we always have $\mathrm{Grad}(\bm{M},l)=0$. This phenomenon is expected due to the specific structure of the attention mechanism. In the last layer, the logit corresponding to the first predicted token is computed solely based on the last query vector and the entire key matrix $\bm{K}^{T}$. Formally, this corresponds to computing only the last row of the attention map, represented as $\mathrm{Softmax}(\bm{Q}_{\mathrm{last}}\bm{K}^{T})$. Since the gradient measure $\mathrm{Grad}(\bm{M},l)$ is defined based on the derivative of the ground truth output logit with respect to parameters associated with the attention sections defined by the mask, and because the first output token logit does not depend on the Middle Q-K sections in the last layer (which involve query vectors from other positions), the resulting gradient naturally becomes zero.
During the training phase, this observation does not apply. The training loss is typically computed using all output logits across the entire input sequence, which requires calculating the full attention matrix in the last layer. However, during inference, since our primary interest lies only in the prediction, we can optimize the computation in the final layer by restricting attention calculation to $\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{Q}_{\mathrm{last}}\bm{K}^{T})$.
Moreover, we observe that the gradient values for the Middle Q-K sections follow a consistent trend across all models: they remain moderate in the shallow layers but drop sharply in deeper layers. As shown in Figure 2, this pattern is evident. For Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, the decline occurs between layers 10 and 20, while for Qwen2.5-7B-Instruct, it starts after layer 20.
This suggests that the attention from the Middle Q-K sections in deep layers contribute less to the output logits. Notably, the number of elements in the Middle Q-K sections scales quadratically with the input token length $N$, whereas the elements in the Streaming and Last Q-K sections grow only linearly with $N$. Therefore, if the computation of the Middle Q-K sections is skipped in deep layers, the time complexity of attention can be reduced from $O(N^{2})$ to $O(N)$. And we assume such skipping will not cause significant loss in output quality.
Based on these observations, we propose TriangleMix, an effective and efficient static attention pattern designed for long-context LLMs. We denote the starting layer of Triangle attention as $L_{\mathrm{tri\_start}}$. For layer $l\leq L_{\mathrm{tri\_start}}$, we apply standard dense attention:
$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M})\right)$
where $\bm{M}$ is the standard causal attention mask.
In layer $l>L_{\mathrm{tri\_start}}$, we skip the Middle Q-K sections and apply a Triangle-shape attention pattern:
$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c\left(1-(\bm{M}-\bm{M}^{\text{middle}})\right)\right)$
where $\bm{M}^{\text{middle}}$ represents the mask corresponding to the Middle Q-K sections.
Thus, in deep layers, the attention complexity becomes $O(N)$, and there is no need for dynamic estimation of the attention mask, making the computation highly efficient. Furthermore, we can optionally adopt dynamic sparse attention in shallow layers to further accelerate computation, while keeping the static Triangle pattern in deep layers.
We acknowledge that the Triangle-shape attention pattern was first introduced in Scbench Li et al. (2024a). However, Scbench applies the Triangle-shape attention uniformly across all layers, leading to notable performance degradation. In contrast, by using a gradient-based probing method, we find that the Middle Q-K sections remain important in shallow layers. Therefore, by combining dense (or dynamic sparse) attention in shallow layers and Triangle-shape attention in deep layers, TriangleMix preserves model performance while achieving significant efficiency gains.
3 Evaluations
3.1 Settings
LLMs. We evaluate our method on the following long-context language models. (1) Llama-3.1-8B-Instruct, a state-of-the-art model from Meta with a context window of up to 128K tokens Grattafiori et al. (2024). (2) Llama-3-8B-Instruct-262K, a fine-tuned variant from Llama-3-8B-Instruct that supports up to 262K tokens GradientAI (2024). (3) Qwen2.5-7B-Instruct, a model from Alibaba with a maximum context window of 128K Yang et al. (2024).
Static Sparsity Baselines. We compare our method with four static sparsity baselines. The first is the Streaming pattern Xiao et al. (2023), applied to all layers with $si=8$ and $sl=512$. The second is Triangle Attention applied to all layers, using $si=8$, $sl=512$, and $last=128$. The third is a hybrid baseline StremingMix, where dense attention is used for layers $l\leq L_{\text{tri\_start}}$, and streaming attention is used for deeper layers.
We also compare against DuoAttention, which learns a separate sparsity pattern for each attention head Xiao et al. (2024). DuoAttention provides a trained pattern for Llama-3.1-8B-Instruct; we train a sparsity pattern for Llama-3-8B-Instruct-262K using the authors’ script. However, for Qwen2.5-7B-Instruct, the training does not converge, so its result is omitted. At the prefilling stage, half of the heads follow the streaming pattern, while the other half use dense attention.
Dynamic Sparsity Baselines. We compare with two dynamic sparsity methods: MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025). We apply default sparsity settings of all LLMs from MInference. For FlexPrefill, we test its hyperparameter $\gamma=0.90$ and $\gamma=0.95$. Additionally, our method can be combined with dynamic sparsity attention: Dynamic sparsity (MInference or FlexPrefill) is applied for layers $l\leq L_{\text{tri\_start}}$, while Triangle Attention is used for deeper layers. This combination reduces runtime overhead because Triangle Attention avoids runtime mask estimation in dynamic sparse attention. We use “Ours + MInference” and “Ours + FlexPrefill” to denote the corresponding combined versions.
Hyperparameters. The hyperparameter $L_{\text{tri\_start}}$ determines the layer from which the Triangle Attention pattern is applied. For layers $l\leq L_{\text{tri\_start}}$, we use dense attention or dynamic sparsity methods. For layers $l>L_{\text{tri\_start}}$, we switch to the static Triangle Attention. Based on the observations in Figure 4, we set $L_{\text{tri\_start}}=16$ for Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, and $L_{\text{tri\_start}}=20$ for Qwen2.5-7B-Instruct. For the Triangle mask, we use: sink tokens $si=8$, sliding window size $sl=512$, and last window size $last=128$, which is the same with the sparse sparsity baselines.
Benchmarks. We evaluate on two benchmarks: (1) RULER: A synthetic long-context benchmark including tasks like hay-in-the-stack, KV retrieval, variable tracking, and QA. We evaluate at input lengths of 4K, 8K, 16K, 32K, 64K, and 128K, with 100 samples per task-length pair Hsieh et al. (2024). (2) LongBench: A realistic long-context benchmark with tasks such as single-doc and multi-doc QA, few-shot learning, code generation, retrieval, summarization, and counting Bai et al. (2023). We evaluate all English tasks.
Specifically, Qwen2.5-7B-Instruct adopts Yarn scaling to improve its long context ability for input length larger than 32K Peng et al. (2023). We report the performance of the origin Qwen2.5-7B-Instruct model for RULER length < 32K, and the Yarn-scaled model for RULER length 32K, 64K and 128K. For LongBench, we only evaluate the origin Qwen2.5-7B-Instruct model for simplicity.
3.2 Effectiveness of TriangleMix
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
Streaming 64.1 55.4 40.5 28.9 26.7 3.3 36.5
Triangle 88.8 88.3 82.8 72.6 65.0 39.0 72.7
StreamingMix 94.3 91.8 90.2 86.2 79.2 71.9 85.6
DuoAttention 95.7 93.1 88.4 84.3 82.5 64.0 84.7
[…]
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
Streaming 49.4 38.7 33.4 30.2 26.5 21.7 33.3
Triangle 88.1 84.7 80.6 71.0 65.0 55.4 74.1
StreamingMix 87.3 83.2 80.8 79.9 77.6 70.8 79.9
DuoAttention 92.8 91.9 89.0 85.0 81.1 76.1 86.0
[…]
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
Streaming 55.6 47.8 35.9 30.6 28.1 21.0 36.5
Triangle 88.4 81.9 74.0 58.3 51.0 14.6 61.4
StreamingMix 93.3 90.4 89.4 81.7 76.4 63.6 82.5
[…]
Comparison with Static Sparsity. Table 3 and Table 2 present the evaluation results of TriangleMix and various static sparsity baselines. TriangleMix consistently outperforms all static baselines and remains nearly lossless performance across all input lengths. In contrast, other static methods usually suffer noticeable performance degradation. DuoAttention performs comparably to TriangleMix on LongBench using Llama-3.1-8B-Instruct. However, its performance is worse than TriangleMix on RULER, especially deteriorates at longer lengths (e.g., 64K and 128K). Furthermore, DuoAttention requires a separate training phase to learn head-wise sparsity patterns, which introduces additional computational overhead and may fail to converge on models such as Qwen-2.5-7B-Instruct.
On the LongBench benchmark, we observe that the full Triangle pattern performs poorly on PassageRetrieval tasks, while the StreamingMix pattern underperforms on the in-context learning task (TREC). These findings suggest that attention over the Middle Q-K section in shallow layers is crucial for retrieval tasks, while attention over the Last Q-K section in deeper layers plays a key role in supporting in-context learning.
Comparison with Dynamic Sparsity. Table 4 and Table 8 report results comparing TriangleMix with dynamic sparsity methods, as well as their combinations. FlexPrefill when $\gamma=0.90$ usually leads to performance degradation, and MInference generally outperforms it. While the performance of MInference and FlexPrefill with $\gamma=0.95$ is similar on the RULER benchmark, MInference still achieves better results on LongBench. When combined with either MInference or FlexPrefill, TriangleMix always fully retains the original performance of the dynamic sparsity methods. Moreover, combining TriangleMix with FlexPrefill even yields slight performance gains, particularly with Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct.
Overall, TriangleMix preserves the full accuracy of dense attention across tasks while significantly reducing time complexity from $O(N^{2})$ to $O(N)$ for layers $l>L_{\text{tri\_start}}$.
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
MInference 96.3 95.1 95.0 90.5 86.8 75.0 89.8
[…]
FlexPrefill (0.90) 90.1 89.7 89.4 89.1 82.0 74.1 85.7
[…]
FlexPrefill (0.95) 95.0 94.7 94.5 92.8 86.5 75.9 89.9
[…]
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
MInference 93.4 90.7 88.8 85.1 82.6 80.1 86.8
[…]
FlexPrefill (0.90) 82.4 80.0 80.7 76.2 74.5 69.1 77.1
[…]
FlexPrefill (0.95) 90.3 87.7 88.0 83.8 80.2 75.7 84.3
[…]
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
MInference 95.6 93.8 92.8 84.9 79.1 68.0 85.7
[…]
FlexPrefill (0.90) 84.4 79.4 82.4 71.4 65.4 47.7 71.8
[…]
FlexPrefill (0.95) 91.1 89.7 88.5 75.4 72.5 51.2 78.1
[…]
Table 4: Comparison with dynamic sparsity methods on RULER.
3.3 Analysis of Start Layer
Figure 5: Average RULER score at 64K length for different $L_{\mathrm{tri\_start}}$ values.
In this section, we investigate the choice of $L_{\text{tri\_start}}$ for each LLM. We evaluate all models on the 64K length tasks in the RULER benchmark, sweeping $L_{\text{tri\_start}}$ from 0 to the total number of layers with a step size of 2. As shown in Figure 5, Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit similar patterns: setting $L_{\text{tri\_start}}=12$ retains 99.7% and 100.0% of their original performance, respectively. This indicates that 62.5% of the layers can adopt the $O(N)$ Triangle Attention without noticeable degradation. In contrast, Qwen-2.5-7B-Instruct requires a higher threshold: setting $L_{\text{tri\_start}}=20$ retains 97.9% of the performance, corresponding to a sparsity ratio of 28.6%.
Method 32K 64K 128K
Dense 44 179 750
MInference 94 (0.5x) 152 (1.2x) 229 (3.3x)
FlexPrefill (0.95) 32 (1.4x) 71  (2.5x) 220 (3.4x)
FlexPrefill (0.90) 24 (1.8x) 52  (3.4x) 143 (5.2x)
[…]
Table 5: Attention kernel latency (ms) for one layer measured on Llama-3.1-8B-Instruct.
Figure 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct.

### v2

@@ 2.4 TriangleMix
Based on these observations, we introduce TriangleMix, an effective and efficient static attention pattern for long-context prefilling. The key idea is that the Middle Q–K section in some layers contributes little to decoding, and can therefore be safely skipped. In these layers, attention reduces to the Triangle pattern shown in Figure 2(b). This modification lowers the complexity of attention in those layers from $O(N^{2})$ to $O(N)$. Unlike dynamic sparsity, Triangle attention is static: there is no need to predict block indices or design specialized sparse kernels. As a result, the kernel is much simpler and faster to implement. Overall, TriangleMix combines dense attention in some layers with Triangle attention in others.
For a given model, the application of TriangleMix is as follows. We first conduct the gradient-based analysis introduced in Sections 2.1 and 2.2. For each layer $l_{i}$, we compute the decoding-time contribution of its Middle Q–K section, denoted as $\mathrm{Grad}(\bm{M}^{\mathrm{middle}},l_{i})$ for $i=1,2,\dots,N_{\mathrm{layer}}$. Based on these values, we identify the $L_{\mathrm{tri}}$ layers with the lowest contributions, where $L_{\mathrm{tri}}$ is a hyperparameter controlling how many layers are converted into Triangle attention. These selected layers adopt the Triangle pattern at inference time, thereby reducing computation. The remaining layers retain their original dense attention. Alternatively, they can also be replaced with dynamic sparse attention methods to achieve further acceleration.
3 Evaluations
3.1 Settings
LLMs and Benchmarks. We evaluate our method on three recent long-context LLMs: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3-8B-Instruct-262K (GradientAI, 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024). We use two challenging benchmarks: LongBench (Bai et al., 2023) and RULER (Hsieh et al., 2024). On RULER, we test input lengths of 4K, 8K, 16K, 32K, 64K, and 128K. On LongBench, we evaluate all English tasks.
Hyperparameters. Different models exhibit varying levels of decoding-time contribution sparsity. We set $L_{\mathrm{tri}}=16$ for Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, and $L_{\mathrm{tri}}=8$ for Qwen2.5-7B-Instruct, which preserves nearly lossless accuracy. An analysis of different $L_{\mathrm{tri}}$ settings is provided in Section 4. For all experiments, we fix the sink token size to 8 and the local window size to 512.
Static Sparsity Baselines. We compare our approach with four static sparsity baselines: (1) Streaming pattern (Xiao et al., 2023; Han et al., 2024), applied to all layers during prefilling; (2) Triangle attention, also applied to all layers; (3) StreamingMix, where Triangle attention in TriangleMix is replaced by the Streaming pattern; (4) DuoAttention (Xiao et al., 2024), which learns a separate sparsity pattern for each attention head. For DuoAttention, we use the official pattern for Llama-3.1-8B-Instruct and train a pattern for Llama-3-8B-Instruct-262K using the authors’ script. We set the sparsity ratio to be 50%. However, training on Qwen2.5-7B-Instruct did not converge, so results are omitted. For all static sparsity baselines, we set the sink token number and local window to 8 and 512, consistent with our method.
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
Streaming 64.1 55.4 40.5 28.9 26.7 3.3 36.5
Triangle 88.8 88.3 82.8 72.6 65.0 39.0 72.7
StreamingMix 94.3 91.8 90.2 86.2 79.2 71.9 85.6
DuoAttention 95.7 93.1 88.4 84.3 82.5 64.0 84.7
TriangleMix 96.3 95.1 94.7 91.3 86.3 77.5 90.2
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
Streaming 49.4 38.7 33.4 30.2 26.5 21.7 33.3
Triangle 88.1 84.7 80.6 71.0 65.0 55.4 74.1
StreamingMix 87.3 83.2 80.8 79.9 77.6 70.8 79.9
DuoAttention 92.8 91.9 89.0 85.0 81.1 76.1 86.0
TriangleMix 93.5 91.0 88.1 85.0 82.4 79.6 86.6
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
Streaming 55.6 47.8 35.9 30.6 28.1 21.0 36.5
Triangle 88.4 81.9 74.0 58.3 51.0 14.6 61.4
StreamingMix 93.3 90.4 89.4 81.7 76.4 63.6 82.5
TriangleMix 95.5 93.8 92.1 84.6 80.2 67.7 85.6
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
TriangleMix 96.3 95.1 94.7 91.3 86.3 77.5 90.2
MInference 96.3 95.1 95.0 90.5 86.8 75.0 89.8
Ours + MInfer 96.2 95.0 94.5 90.8 87.2 75.8 89.9
FlexPrefill 95.0 94.7 94.5 92.8 86.5 75.9 89.9
Ours + FP 95.2 95.0 94.7 92.8 86.8 76.2 90.1
XAttn 96.5 94.9 94.5 91.8 85.7 73.8 89.5
Ours + XAttn 96.2 95.1 94.9 91.9 85.7 72.9 89.4
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
TriangleMix 93.5 91.0 88.1 85.0 82.4 79.6 86.6
MInference 93.4 90.7 88.8 85.1 82.6 80.1 86.8
Ours + MInfer 93.0 90.4 88.6 84.4 82.7 80.1 86.5
FlexPrefill 90.3 87.7 88.0 83.8 80.2 75.7 84.3
Ours + FP 90.2 87.3 87.3 83.8 80.0 76.6 84.2
XAttn 93.6 91.1 87.7 84.8 82.6 78.9 86.5
Ours + XAttn 93.3 91.2 87.9 84.3 82.5 79.3 86.4
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
TriangleMix 95.5 93.8 92.1 84.6 80.2 67.7 85.6
MInference 95.6 93.8 92.8 84.9 79.1 68.0 85.7
Ours + MInfer 95.7 93.5 92.4 84.8 76.9 63.5 84.5
FlexPrefill 91.1 89.7 88.5 75.4 72.5 51.2 78.1
Ours + FP 92.2 91.1 89.5 77.0 73.3 52.7 79.3
XAttn 95.4 93.7 92.0 82.9 77.8 66.2 84.7
Ours + XAttn 95.3 93.3 91.5 82.2 77.4 64.9 84.1
Table 3: Comparison with dynamic sparsity methods on RULER.
Dynamic Sparsity Baselines. We further compare with three dynamic sparsity methods: MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025). Our method can also be combined with these approaches: Triangle attention is retained, while layers that require full attention are replaced with dynamic sparse attention. This design further reduces runtime overhead compared with dynamic sparsity alone, since Triangle attention skips more attention computation and is much simpler. We denote the combined versions as “Ours + MInference,” “Ours + FlexPrefill,” and “Ours + XAttention.” For dynamic methods, we also set their hyperparameters to maintain near-lossless accuracy: $\gamma=0.95$ for FlexPrefill, and $\tau=0.95$ with $\mathrm{stride}=8$ for XAttention.
3.2 Effectiveness of TriangleMix
Comparison with Static Sparsity. Table 3 and Table 5 present the evaluation results of TriangleMix and various static sparsity baselines. TriangleMix consistently outperforms all static baselines and remains nearly lossless performance across all input lengths. In contrast, other static methods usually suffer noticeable performance degradation. DuoAttention performs comparably to TriangleMix on LongBench using Llama-3.1-8B-Instruct. However, its performance is worse than TriangleMix on RULER, especially deteriorates at longer lengths (e.g., 64K and 128K). Furthermore, DuoAttention requires a separate training phase to learn head-wise sparsity, which introduces additional computational overhead and may fail to converge on models such as Qwen-2.5-7B-Instruct.
On the LongBench benchmark, we observe that the full Triangle pattern performs poorly on PassageRetrieval tasks, while the StreamingMix pattern underperforms on the in-context learning task (TREC). These findings suggest that attention over the Middle Q-K section in certain layers (especially shallow layers) is still crucial for retrieval tasks, while attention over the Last Q-K section in certain layers (especially deep layers) plays a key role in supporting in-context learning.
Methods Average
2Wiki
GovRep
Hotpot
LCC
MNews
MF-en
PsgCnt
PsgRtr
Qasper
Rbench
Samsum
TREC
TrvQA
TriangleMix 54.5 47.8 34.3 61.5 67.0 25.8 56.5 16.8 98.3 44.8 51.9 42.2 69.7 92.2
Ours + MInference 54.5 48.3 34.2 62.0 66.7 25.6 56.6 17.4 98.3 44.8 51.6 42.1 69.3 92.1
FlexPrefill 48.4 39.4 33.7 58.3 67.2 25.6 55.6 4.0 47.7 42.5 53.8 41.9 69.7 90.2
Ours + FlexPrefill 52.2 39.4 33.5 58.5 66.9 25.7 54.5 3.3 97.0 43.1 53.8 42.3 69.3 90.8
XAttention 53.5 45.4 34.5 61.0 66.9 25.7 55.9 18.1 86.0 42.9 55.5 42.9 71.0 90.2
Ours + XAttention 54.3 45.0 34.6 61.0 66.8 25.6 57.1 19.0 98.7 42.4 54.4 42.0 69.3 90.4
TriangleMix 43.4 20.7 34.5 25.7 47.3 26.1 51.1 4.7 85.3 30.7 43.6 40.9 66.7 87.0
Ours + MInference 43.4 20.6 34.4 26.4 47.2 26.2 51.1 4.3 85.3 30.3 43.6 40.9 66.7 86.9
FlexPrefill 35.0 17.7 33.0 25.1 36.8 26.1 49.1 3.3 13.7 28.3 36.2 39.1 65.0 81.7
Ours + FlexPrefill 36.0 18.3 33.2 25.1 36.9 26.2 49.6 6.7 24.0 28.3 35.6 38.9 64.3 80.4
XAttention 41.5 20.7 34.3 26.0 37.9 26.1 52.3 4.7 71.3 30.8 39.4 40.0 68.3 87.2
Ours + XAttention 42.1 21.1 34.3 26.1 37.7 26.2 51.7 5.7 81.3 30.3 38.3 39.9 67.3 87.0
TriangleMix 46.1 24.7 30.6 32.3 62.6 22.3 40.9 13.1 89.3 23.9 59.6 42.4 69.3 88.5
Ours + MInference 46.1 25.7 30.3 32.2 61.9 22.3 40.7 11.4 89.7 25.8 59.4 42.5 68.7 89.0
FlexPrefill 40.8 20.2 29.7 33.9 66.8 21.8 40.4 4.0 48.7 17.1 54.8 40.5 67.3 85.5
Ours + FlexPrefill 41.7 21.0 29.8 32.8 64.8 21.9 39.6 5.3 56.0 19.8 55.1 40.9 69.0 85.8
XAttention 46.6 22.4 30.9 34.0 63.9 22.4 40.5 10.9 98.0 23.5 60.9 42.2 67.7 88.8
Ours + XAttention 45.8 23.6 31.1 33.1 63.8 21.7 41.3 9.8 88.3 24.5 58.2 42.4 69.0 88.4
Table 4: Comparison with dynamic sparsity methods on LongBench.
Comparison with Dynamic Sparsity. Tables 3 and 4 present the results of TriangleMix compared with dynamic sparsity methods, as well as their combinations. Despite its simplicity, TriangleMix maintains accuracy comparable to dynamic sparsity methods. On the RULER benchmark, it yields an average accuracy change of only $0.4\%$ to $0.1\%$ relative to full attention, which is on par with MInference ($0.6\%$ to $0.3\%$), and outperforms FlexPrefill ($9.1\%$ to $0.6\%$) and XAttention ($2.2\%$ to $0.0\%$). Importantly, although TriangleMix skips many attention blocks with non-trivial scores, the decoding accuracy remains nearly lossless. This supports our claim that such blocks contribute only marginally to decoding-time accuracy, and that decoding-time sparsity is a general property across all three tested models.
Moreover, when combined with dynamic sparsity methods such as MInference, FlexPrefill, or XAttention, TriangleMix consistently preserves their original performance. On LongBench, combining TriangleMix with FlexPrefill or XAttention even yields slight performance improvements. These combinations are meaningful: by eliminating computation in many additional attention blocks, TriangleMix reduces the computation required by dynamic sparsity methods, making prefilling even faster than using dynamic sparsity alone.
Overall, TriangleMix preserves nearly full accuracy of dense attention across tasks while significantly reducing time complexity from $O(N^{2})$ to $O(N)$ in layers using Triangle attention.
@@ 3.2 Effectiveness of TriangleMix
TriangleMix 46.1 24.7 30.6 32.3 62.6 22.3 40.9 13.1 89.3 23.9 59.6 42.4 69.3 88.5
Table 5: Comparison with static sparsity methods on Longbench.
3.3 Analysis of $L_{\mathrm{tri}}$
Figure 4: Average RULER score at 64K length for different $L_{\mathrm{tri}}$ values.
In this section, we investigate the choice of $L_{\text{tri}}$ for each LLM. We evaluate all models on the 64K length tasks in the RULER benchmark, sweeping $L_{\text{tri}}$ from 0 to the total number of layers with a step size of 2. As shown in Figure 4, Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit similar patterns: setting $L_{\text{tri}}=20$ retains 99.7% and 100.0% of their original performance, respectively. This indicates that 62.5% of the layers can adopt the $O(N)$ Triangle Attention without noticeable degradation. In contrast, Qwen-2.5-7B-Instruct requires a less threshold: setting $L_{\text{tri}}=8$ retains 98.2% of the performance, corresponding to a sparsity ratio of 28.6%.
Llama-3.1-8B-Instruct Qwen2.5-7B-Instruct
Method 32K 64K 128K 32K 64K 128K
Dense 44 179 750 40 156 646
MInference 94 (0.5x) 152 (1.2x) 229 (3.3x) 93 (0.4x) 151 (1.0x) 245 (2.6x)
FlexPrefill 32 (1.4x) 71 (2.5x) 220 (3.4x) 38 (1.0x) 99 (1.6x) 310 (2.1x)
XAttention 47 (0.9x) 127 (1.4x) 391 (1.9x) 44 (0.9x) 128 (1.2x) 391 (1.7x)
Triangle 12 (3.7x) 24 (7.5x) 49 (15.3x) 11 (3.7x) 21 (7.4x) 76 (8.5x)
Table 6: Attention kernel latency (ms) for one layer.